# Coexistence of high electron-mobility, unpaired spins, and superconductivity at high carrier density SrTiO3-based interfaces

### v1

@@ Abstract
The t2g band-structure of SrTiO3-based two-dimensional electron gasses (2DEGs), have been found to play a role in features such as the superconducting dome, high-mobility transport, and the magnitude of spin-orbit coupling. This adds to the already very diverse range of phenomena, including magnetism and extreme magnetoresistance, exhibited by this particular material platform. Tuning and/or combining these intriguing attributes could yield significant progress within quantum and spintronics technologies. Doing so demands precise control of the parameters, which requires a better understanding of the factors that affect them. Here we present effects of the t2g band-order inversion, stemming from the growth of spinel-structured -Al2O3 onto perovskite SrTiO3. Electronic transport measurements show that with LaAlO3/SrTiO3 as the reference, the carrier density and electron mobility are enhanced, and the sample displays a reshaping of the superconducting dome. Additionally, unpaired spins are evidenced by increasing Anomalous Hall Effect with decreasing temperature, entering the same temperature range as the superconducting transition, and a Kondo-like upturn in the sheet resistance. Finally, it is argued that the high-mobility dxz/yz-band is more likely than the dxy-band to host the supercurrent.
@@ I Introduction
Furthermore, the relation between charge carrier density, ns, mobility, and superconducting critical temperature, TC, of -Al2O3/SrTiO3 is discussed and compared to previous studies. Based on the two -Al2O3/SrTiO3 samples measured in this study, we arrive at indications of a novel reshaping of the superconducting dome with the changed band-ordering. Taking Anomalous Hall Effect (AHE) and Kondo-like features in the electronic transport data into account indicates the coexistence of unpaired spins and superconductivity in the same temperature range. Finally, we discuss the origin of the two-dimensional superconducting phase, whether this arises from the dxy-band or dxz/yz-band, and how this ties into the reshaped superconducting dome.
@@ II Experimental
Six samples were investigated in this study. Sample 1, 2, 3 and 6 are -Al2O3( 5 nm)/SrTiO3(0.5 mm) grown by Pulsed Laser Deposition (PLD) for 4 min. Sample 1 and 2 was grown with a laser fluence of 9 J/cm2 and an oxygen partial pressure at 2 10-6 mbar, sample 3 was grown with a laser fluence of 6 J/cm2 at an oxygen partial pressure at 2 10-6 mbar. Sample 6 was grown with a laser fluence of 9 J/cm2 at an oxygen partial pressure at 1 10-6 mbar, but with a cooldown pressure after deposition at 1 10-7 mbar. Sample 4 is a -Al2O3( 34.7 nm)/SrTiO3(0.5 mm) grown with a laser fluence of 9 J/cm2 at an oxygen partial pressure at 1 10-6 mbar for 30 min, and sample 5 is a LaAlO3/SrTiO3 sample grown by Z. Liao at the University of Twente (Twente, The Netherlands) further described in ref. [34].
@@
Measurement System equipped with a 16 T magnet, and below 2 K, the electronic measurements were performed in a dilution refrigerator with a 6-1-1 T vector magnet.
@@ III Structural comparison
HRSTEM images of LaAlO3/SrTiO3 (sample 5) and -Al2O3/SrTiO3 (sample 3) are displayed in Fig. 1a) and d), respectively. These show a atomic ordering corresponding to epitaxial growth. We expect this to be representative for the entire sample consistent with the Kiessig fringes visible in XRR of Sample 4 in Supplemental Material (SM) III. As illustrated in Fig. 1b), the interface between LaAlO3 and SrTiO3 consists of two cubic perovskites with a lattice mismatch of 3.0%. The comparatively lower lattice constant of LaAlO3 induces compressive strain in SrTiO3 near the interface.
@@
DFT calculated band structure of these interfaces are presented in Fig. 1c) for LaAlO3/SrTiO3 [3] and Fig. 1f) for -Al2O3/SrTiO3 [4]. Here we observe an inversion of the dxy and dxz/yz bands in the range of the Fermi energy consistent with data from resonant soft x-ray linear dichroism[27] and angle-resolved photoemission spectroscopy.[4]
@@ IV Low-Temperature Transport
In Fig. 2b) the out-of-plane magnetoresistance, MR , is shown with i = a, b, or s for the two vdP configurations, and the combined sheet resistance respectively. Worth noting is the resistance anisotropy between the vdP configurations, which increase with magnetic field such that . The overall magnetoresistance reaches the range of 103% at 6 T in both vdP configurations and does not saturate, placing this sample in the extreme magnetoresistance regime.[38]. Worth noticing is the absence of features indicating weak localization (WL), which otherwise could have provided an alternative explanation to the resistance upturn below 1.5 K in Fig. 2a).[39]
@@
In Fig. 3d), is normalized with Rs, but shows an increase with decreasing temperature. This is consistent with the 50 mK Hall plot displayed in Fig 3a), which has the most prominent AHE indicative slope of all the data. The seems to be temperature independent, fluctuating below 6 T throughout the temperatures.
@@
The observation of AHE can be explained by scattering from unpaired spins and is consistent with the Kondo-like Rs upturn below 1.5 K in Fig. 2a). The upturn could also result from freeze-out of nano- or even microscale regions of the 2DEG.[8, 9, 49, 50, 51] Since the energy density of the PLD plume is highest at the centre of the plume front, where the species have travelled the least when they reach the substrate, it can be expected that the 2DEG is inhomogeneous. In addition SrTiO3 furthermore undergoes a ferroelastic transition from cubic to tetragonal at 105 K, creating domains of different tetragonal orientations.[52] Regarding freeze-out, this is not directly consistent with the AHE signature, leaving the Kondo-like explanation as the most likely.
@@ V Superconductivity
The temperature dependence of the differential resistance, , is shown in Fig. 4b, with the superconductivity-related non-linearities suppressed as approaches TC. To estimate the critical temperature, , we extract in the range nA, as indicated by the dashed lines in Fig 4b), and display the average of this as a function of temperature in Fig. 4c). The gradual transition between the superconducting and normal state is typical for STO-based 2DEGs.
@@ VI Discussion
Considering the band ordering of the dxy-band compared to the dxz/yz-band, we can tentatively assign the superconductivity as arising from the dxz/yz-band. If the superconductivity came from the dxy-band, the QCP would mark the onset of filling this band. Increasing ns would increase the critical values until a maximum (TCmax) marking the start of the filling of the subsequent dxz/yz-band. This is inconsistent with the band inversion of -Al2O3/SrTiO3, which shows superconductivity, while the dxz/yz-band is lower in energy than the dxy-band and therefore strongly dominating. Contrarily, if the superconductivity came from dxz/yz-band, the QCP would witness the onset of filling into this band and the TCmax the filling into the dxy-band. This is consistent with the superconductivity at a broader range and even higher carrier densities than in LaAlO3/SrTiO3.
One further detail worth noting is the increasing AHE contribution at low temperatures, which keeps increasing in Sample 1 down to 50 mK below which Hall data were not measured. This temperature, however, overlaps with the Gausian distribution describing the critical temperatures of the puddles, suggesting the co-existence of superconductivity and unpaired spins in the same temperature range in Sample 1. This is further supported by the upturn in Rs at the lowest temperatures until the superconducting transition. As previously described in ref. [41], these unpaired spins likely originate from the oxygen vacancies existing at and near the interface in -Al2O3/SrTiO3. This is consistent with the understanding that the 2DEG is generated by electrons donated by oxygen vacancies, and they should not disappear as superconductivity arises but indeed result in an effective lowering of TC compared to LaAlO3/SrTiO3 as indicated by Fig. 5. Additionally, because one oxygen vacancy donates one electron, the oxygen vacancy density should be on the same order of magnitude as the carrier density of the electrons.
@@ VII Conclusion
Below T 100 mK, the -Al2O3/SrTiO3 interfaces showed superconductivity consistent with a distribution of Josephson-Junction coupled superconducting puddles in the interface. The order-of-magnitude larger carrier density at which superconductivity was observed compared to LaAlO3/SrTiO3 suggests that the band inversion significantly reshapes the superconducting dome. Furthermore, due to the presence of superconductivity in the strongly dxz/yz-band dominated interface, it is unlikely that the dxy-band is responsible for SC transport in -Al2O3/SrTiO3. Finally, we see indications of unpaired spins in one of the -Al2O3/SrTiO3 samples through Anomalous Hall Effect and Kondo-like features. The Anomalous Hall Effect resistance contribution tends to increase with decreasing temperature, yielding the strongest signature at temperatures that overlaps with the Gaussian distribution of superconducting puddle TCs. Thus the existence of superconductivity and unpaired spins, likely stemming from oxygen vacancies, in the same temperature range seems to be a property of -Al2O3/SrTiO3.
@@ VIII Acknowledgements
We thank V. Rosendal, C. E. N. Petersen, R. T. Dahm, H. Witt, S. Mallik, and M. Bibes for helpful discussions. We thank Z.L. Liao and G. Koster at the University of Twente for providing the LAO/STO sample used to compare STEM images. T.H.O. and F.T. acknowledge support by research grant 37338 (SANSIT) from Villum Fonden. T.S.J. acknowledges support from the Novo Nordic Foundation Challenge Program, grant no. NNF21OC0066526 (BioMag). N.G. and J.V. acknowledge funding from GOA project “Grignard 2.0” of the University of Antwerp. Z.Zhang is gratefuly acknowledged for image processing of the STEM image stacks.

### v2

@@ Abstract
The t2g band-structure of SrTiO3-based two-dimensional electron gasses (2DEGs), have been found to play a role in features such as the superconducting dome, high-mobility transport, and the magnitude of spin-orbit coupling. This adds to the already very diverse range of phenomena, including magnetism and extreme magnetoresistance, exhibited by this particular material platform. Tuning and/or combining these intriguing attributes could yield significant progress within quantum and spintronics technologies. Doing so demands precise control of the parameters, which requires a better understanding of the factors that affect them. Here we present effects of the t2g band-order inversion, stemming from the growth of spinel-structured -Al2O3 onto perovskite SrTiO3. Electronic transport measurements show that with LaAlO3/SrTiO3 as the reference, the carrier density and electron mobility are enhanced, and the sample displays a reshaping of the superconducting dome. Additionally, unpaired spins are evidenced by increasing Anomalous Hall Effect with decreasing temperature, entering the same temperature range as the superconducting transition. Finally, it is argued that the high-mobility dxz/yz-band is more likely than the dxy-band to host the supercurrent.
@@ I Introduction
Furthermore, the relation between charge carrier density, ns, mobility, and superconducting critical temperature, TC, of -Al2O3/SrTiO3 is discussed and compared to previous studies. Based on the two -Al2O3/SrTiO3 samples measured in this study, we arrive at indications of a novel reshaping of the superconducting dome with the changed band-ordering. Taking Anomalous Hall Effect (AHE) in the electronic transport data into account indicates the coexistence of unpaired spins and superconductivity in the same temperature range. Finally, we discuss the origin of the two-dimensional superconducting phase, whether this arises from the dxy-band or dxz/yz-band, and how this ties into the reshaped superconducting dome.
@@ II Experimental
Six samples were investigated in this study. Sample 1, 2, 3,and 6 are -Al2O3( 6-7 nm)/SrTiO3(0.5 mm) grown by Pulsed Laser Deposition (PLD) for 4 min. Sample 1 and 2 was grown with a laser fluence of 9 J/cm2 and an oxygen partial pressure at 2 10-6 mbar, sample 3 was grown with a laser fluence of 6 J/cm2 at an oxygen partial pressure at 1 10-6 mbar. Sample 6 was grown with a laser fluence of 9 J/cm2 at an oxygen partial pressure at 1 10-6 mbar, but with a cooldown pressure after deposition at 1 10-7 mbar. Sample 4 is a -Al2O3( 34.7 nm)/SrTiO3(0.5 mm) grown with a laser fluence of 9 J/cm2 at an oxygen partial pressure at 1 10-6 mbar for 30 min, and sample 5 is a LaAlO3/SrTiO3 sample grown by Z. Liao at the University of Twente (Twente, The Netherlands) further described in ref. [34]. The actual fluence value was determined at each deposition based on regular measurements of the viewport transmission, laser energy before entering the PLD chamber as well as spot size measurements carried out on a silver target.
@@
Measurement System equipped with a 16 T magnet, and below 2 K, the electronic measurements were performed in a dilution refrigerator with a 6-1-1 T vector magnet. Due to the low resistances, stemming from the high mobilities and carrier densities, the measurements, showing normal metal conduction, were performed using bias currents on A-scale. In contrast, the small superconducting critical currents led to the measurements of superconductivity using bias currents of magnitudes 10-100 nA.
@@ III Structural comparison
HRSTEM images of LaAlO3/SrTiO3 (sample 5) and -Al2O3/SrTiO3 (sample 3) are displayed in Fig. 1a) and d), respectively. These show a atomic ordering corresponding to epitaxial growth. We expect this to be representative for the entire sample consistent with the Kiessig fringes visible in XRR of Sample 4 in Supplemental Material (SM) III. We note that the growth rate extracted from the Kiessig fringes in the XRR, returning a 5 nm thickness for the samples 1-3 and 6, only provides a lower limit. This is because of the continuous drop in fluence from decreasing viewport transmission during the extended growth time of the thick films. The thickness of sample 4 was calculated manually, as the fitting curves returned inconsistent results. More information can be found in SM III. As illustrated in Fig. 1b), the interface between LaAlO3 and SrTiO3 consists of two cubic perovskites with a lattice mismatch of 3.0%. The comparatively lower lattice constant of LaAlO3 induces compressive strain in SrTiO3 near the interface.
@@
DFT calculated band structure of these interfaces are presented in Fig. 1c) for LaAlO3/SrTiO3 [3] and Fig. 1f) for -Al2O3/SrTiO3 [4]. Here we observe an inversion of the dxy and dxz/yz bands in the range of the Fermi energy consistent with data from resonant soft x-ray linear dichroism[27] and angle-resolved photoemission spectroscopy.[4] One should note that the slopes of the bands in Fig. 1f) do not fully reflect the ones measured by resonant soft-X-ray angle-resolved photoelectron spectroscopy in ref. 4. Hence, Fig. 1f) only indicates the band-order in E.
@@ IV Low-Temperature Transport
In Fig. 2b) the out-of-plane magnetoresistance, MR , is shown with i = a, b, or s for the two vdP configurations, and the combined sheet resistance respectively. Worth noting is the resistance anisotropy between the vdP configurations, which increase with magnetic field such that . To our knowledge, such anisotropy has not previously been reported in -Al2O3/SrTiO3. The overall magnetoresistance reaches the range of 103% at 6 T in both vdP configurations and does not saturate, placing this sample in the extreme magnetoresistance regime.[38]. Worth noticing is the absence of features indicating weak localization (WL), which otherwise could have provided an alternative explanation to the resistance upturn below 1.5 K in Fig. 2a).[39]
@@
. b) Differential resistance as a function of source-drain current and temperature. c) Average dV/dI as a function of temperature, corresponding to the local resistance in b) for Isd between 0.08 A, as indicated by the dashed lines in b). The line in c) is the EMT fit. The uncertainties of each fitting parameter correspond to the change in the parameter necessary to increase with reduced as described in SM IV. d) Differential resistance as a function of source-drain current and out-of-plane magnetic field magnitude.
@@
In Fig. 3d), is normalized with Rs, but shows an increase with decreasing temperature. This is consistent with the 50 mK Hall plot displayed in Fig 3a), which has the most prominent AHE indicative slope of all the data. However, the large increase in the contribution from AHE with decreasing temperature is likely an effect of the diminishing resistance that appears from the superconducting transition in addition to a finite increase in AHE with decreasing temperature as the unpaired spins becomes more dominant. The seems to be temperature independent, fluctuating below 6 T throughout the temperatures.
@@
The observation of AHE can be explained by scattering from unpaired spins and is consistent with the Rs upturn below 1.5 K in Fig. 2a). The upturn could also result from freeze-out of nano- or even microscale regions of the 2DEG.[8, 9, 45, 46, 47] Since the energy density of the PLD plume is highest at the centre of the plume front, where the species have travelled the least when they reach the substrate, it can be expected that the 2DEG is inhomogeneous. In addition SrTiO3 furthermore undergoes a ferroelastic transition from cubic to tetragonal at 105 K, creating domains of different tetragonal orientations.[48] Hence freeze-out and/or a spin-related scattering mechanism could play a role in the low-temperature resistance upturn.
Especially for sample 2 (See SM V), the 2D nature of the conducting interface is dubious. However, recent work from Rubi et al.[49] shows that other oxide 2DEGs (based on SrTiO3 or KTaO2) conduct through a superposition of 2D and 3D states. Additionally, the mobilities in both samples shown here are higher or on the scale of the record mobility in bulk SrTiO3.[50] Hence, it seems unlikely that the high-mobility transport occurs through 3D rather than 2D states, as the latter tends to have higher mobility.
@@ V Superconductivity
The temperature dependence of the differential resistance, , is shown in Fig. 4b, with the superconductivity-related non-linearities suppressed as approaches TC. To estimate the critical temperature, , we extract in the range nA, as indicated by the dashed lines in Fig 4b), and display the average of this as a function of temperature in Fig. 4c). The gradual transition between the superconducting and normal state is typical for SrTiO3-based 2DEGs.
@@ VI Discussion
Considering the band ordering of the dxy-band compared to the dxz/yz-band, we can tentatively assign the superconductivity as arising from the dxz/yz-band. If the superconductivity came from the dxy-band, the QCP would mark the onset of filling this band. Increasing ns would increase the critical values until a maximum (TCmax) marking the start of the filling of the subsequent dxz/yz-band. This is inconsistent with the band inversion of -Al2O3/SrTiO3, which shows superconductivity, while the dxz/yz-band is lower in energy than the dxy-band and therefore strongly dominating. Contrarily, if the superconductivity came from dxz/yz-band, the QCP would witness the onset of filling into this band and the TCmax the filling into the dxy-band. This case was also covered for LaAlO3/SrTiO3 by Maniv et al.[57] presenting a gate dependent non-monotonic dxz/yz band occupation responsible for corresponding non-monotonic Shubnikov-de Haas oscillations and . A superconducting dxz/yz band in -Al2O3/SrTiO3 is consistent with the superconductivity at a broader range and even higher carrier densities than in LaAlO3/SrTiO3. We note that according to Chikina et al.[4] Sample 1, based on its carrier density and high mobility, should be completely characterized by the dxz/yz-band at the Fermi energy. This emphasizes the dxz/yz-band as the host of superconductivity and the reshaped superconducting dome, since the dome-shape is explained by a competition between the dxz/yz-band and dxy-band. Hence, the difference between the superconducting phase-diagrams of LaAlO3/SrTiO3 and -Al2O3/SrTiO3, may originate from their vastly different band structures around the Fermi energy.
One further detail worth noting is the increasing AHE contribution at low temperatures, which keeps increasing in Sample 1 down to 50 mK below which Hall data were not measured. This temperature, however, overlaps with the Gausian distribution describing the critical temperatures of the puddles, suggesting the co-existence of superconductivity and unpaired spins in the same temperature range in Sample 1. This is further supported by the upturn in Rs at the lowest temperatures until the superconducting transition. As previously described in ref. [41], these unpaired spins likely originate from the oxygen vacancies existing at and near the interface in -Al2O3/SrTiO3. This is consistent with the understanding that the 2DEG is generated by electrons donated by oxygen vacancies, and they should not disappear as superconductivity arises but indeed result in an effective lowering of TC compared to LaAlO3/SrTiO3 as indicated by Fig. 5. Additionally, because one oxygen vacancy donates one electron, the oxygen vacancy density should be on the same order of magnitude as the carrier density of the electrons. We note that the oxygen vacancies themselves might not necessarily carry magnetic moments of significant magnitude but could stabilize the cationic defects that do.[58, 41]
@@ VII Conclusion
Below T 100 mK, the -Al2O3/SrTiO3 interfaces showed superconductivity consistent with a distribution of Josephson-Junction coupled superconducting puddles in the interface. The order-of-magnitude larger carrier density at which superconductivity was observed compared to LaAlO3/SrTiO3 suggests that the band inversion significantly reshapes the superconducting dome. Furthermore, due to the presence of superconductivity in the strongly dxz/yz-band dominated interface, it is unlikely that the dxy-band is responsible for superconducting transport in -Al2O3/SrTiO3. Finally, we see indications of unpaired spins in one of the -Al2O3/SrTiO3 samples through Anomalous Hall Effect. The Anomalous Hall Effect resistance contribution tends to increase with decreasing temperature, yielding the strongest signature at temperatures that overlaps with the Gaussian distribution of superconducting puddle TCs. Thus the existence of superconductivity and unpaired spins, likely stemming from oxygen vacancies, in the same temperature range seems to be a property of -Al2O3/SrTiO3.
To the best of our knowledge, neither anisotropy of the out-of-plane magnetoresistance nor superconductivity co-existing with unpaired spins has been reported previously in -Al2O3/SrTiO3. Additionally, the combination of a high electron mobility and a superconducting regime in a single oxide interface is novel.
@@ VIII Acknowledgements
We thank V. Rosendal, C. E. N. Petersen, R. T. Dahm, H. Witt, S. Mallik, and M. Bibes for helpful discussions. We thank Z.L. Liao and G. Koster at the University of Twente for providing the LAO/STO sample used to compare STEM images. T.H.O. and F.T. acknowledge support by research grant 37338 (SANSIT) from Villum Fonden. T.S.J. acknowledges support from the Novo Nordic Foundation Challenge Program, grant no. NNF21OC0066526 (BioMag). N. B. acknowledges support from the ANR Project SURIKAT. N.G. and J.V. acknowledge funding from GOA project “Grignard 2.0” of the University of Antwerp. Z.Zhang is gratefuly acknowledged for image processing of the STEM image stacks.